# On the kernels of the pro-$p$ outer Galois representations associated to once-punctured CM elliptic curves

### v2

@@ 4.1. Construction of elements
In this subsection, we construct elements $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ satisfying the assumption of Conjecture 2.8 in such a way that can be regarded as a two-variable variant of Sharifi’s construction in [Sha02, 2]. We also construct auxiliary elements $\{g_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$, which are similar to $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ but more easier to handle group-theoretically. These auxiliary elements play an important role when proving that $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ strongly generate $F^{1}G$ (cf. the proof of Theorem 2.12 in the next subsection)
First, we lift generators of $\mathrm{Gal}(K(p^{\infty})/K)\cong\Delta\times\mathbb{Z}_{p}^{2}$ to $\mathrm{Gal}(\Omega/K)$, as follows: Let us denote the maximal pro-$p$ subextension of $K(p^{\infty})/K$ by $K_{\infty}/K$. The upper exact sequence in the following commutative diagram
@@
[…]
splits since $\mathrm{Gal}(\Omega/K(p^{\infty}))$ is a pro-$p$ group and $\Delta$ is a prime-to-$p$ group by Schur-Zassenhaus [Ser97, 5.9, Corollary 1]. We fix a section $r\colon\mathrm{Gal}(K(p)/K)\to\mathrm{Gal}(\Omega/K_{\infty})$ and identify $\mathrm{Gal}(\Omega/K)$ with the semi-direct product $G\rtimes\Delta$. Then the group $\Delta$ acts on $\mathrm{Gal}(\Omega/K(p^{\infty}))$ through this section by conjugation, and
$1\to\mathrm{Gal}(\Omega/K(p^{\infty}))\to G\to\mathrm{Gal}(K(p^{\infty})/K(p))\to 1$
is an exact sequence of pro-$p$ groups with $\Delta$-action, noting that $\Delta$ acts trivially on $\mathrm{Gal}(K(p^{\infty})/K(p))$. Then the sequence
$1\to\mathrm{Gal}(\Omega/K(p^{\infty}))^{\Delta}\to G^{\Delta}\to\mathrm{Gal}(K(p^{\infty})/K(p))\to 1$
@@
$G^{\Delta}\to\mathrm{Gal}(K(p^{\infty})/K(p))$
is surjective by applying Lemma 4.1 in the case of $m=0$ below.
Let $\gamma_{1}$ (resp. $\gamma_{2}$) be an element of the $\Delta$-invariant subgroup $G^{\Delta}$ which restricts to a generator of $\mathrm{Gal}(K(p^{\infty})/K(\bar{\mathfrak{p}}^{\infty}\mathfrak{p}))\cong\mathbb{Z}_{p}$ (resp. $\mathrm{Gal}(K(p^{\infty})/K(\mathfrak{p}^{\infty}\bar{\mathfrak{p}}))\cong\mathbb{Z}_{p}$). Moreover, we fix a generator $\delta\in\mathbb{F}_{p}^{\times}$ and let $\delta_{1}$ (resp. $\delta_{2}$) be the image of $(\delta,1)$ (resp. $(1,\delta)$) in $\Delta$ through the isomorphism
$\Delta\xrightarrow[(\chi_{1},\chi_{2})]{\sim}(\mathbb{F}_{p}^{\times})^{\oplus 2}/\mathrm{im}(O_{K}^{\times})$
under the section $r\colon\Delta\to\mathrm{Gal}(\Omega/K)$. Then the element $\delta_{1}$ (resp. $\delta_{2}$) is contained in $\mathrm{Gal}(\Omega/K(\bar{\mathfrak{p}}))$ (resp. $\mathrm{Gal}(\Omega/K(\bar{p}))$) and, by construction, we have the following relations:
$[\delta_{1},\delta_{2}]=1,\quad[\delta_{1},\gamma_{1}]=1,\quad[\delta_{1},\gamma_{2}]=1,\quad[\delta_{2},\gamma_{1}]=1\quad\text{and}\quad[\delta_{2},\gamma_{2}]=1.\quad$
For $m\in\mathbb{Z}_{\geq 0}$ and $i=1,2$, we define $\epsilon_{i,m}\in\mathbb{Z}_{p}[\mathrm{Gal}(\Omega/K)]$ to be
@@
For every $g\in G$, $m\in\mathbb{Z}_{\geq 0}$ and $i=1,2$, the limit
@@
Let $A$ denote the maximal abelian quotient $\mathrm{Gal}(\Omega/K(p^{\infty}))^{{\rm ab}}$ of $F^{1}G=\mathrm{Gal}(\Omega/K(p^{\infty}))$. Then $A$ is naturally endowed with the structure of a $\Lambda\coloneqq\mathbb{Z}_{p}[[\mathrm{Gal}(K(p^{\infty})/K(p))]]$-module. In the following lemma and remark, we determine the structure of this $\Lambda$-module.
@@
Assume that the class number of $K(p)$ is not divisible by $p$, and there are only two primes of $K(p^{\infty})$ above $p$. Let $\boldsymbol{m}\in I_{0}$ and $A^{\boldsymbol{m}}\coloneqq\epsilon_{1,m_{1}}\epsilon_{2,m_{2}}A$. Then the following assertions hold:
@@
    $A^{(p,1)}=A^{(1,p)}$ is a quotient of the annihilator ideal $\mathrm{ann}_{\Lambda}(\mathbb{Z}_{p}(1))$ of the Tate twist $\mathbb{Z}_{p}(1)$, i.e. a quotient of the $\Lambda$-module $M$ generated by two elements $v_{1},v_{2}$ satisfying $(\gamma_{1}-\chi_{1}(\gamma_{1}))v_{2}=(\gamma_{2}-\chi_{2}(\gamma_{2}))v_{1}$.
@@
00 0 $\to H^{1}(\mathrm{Gal}(K(p^{\infty})/K(p)),\mathbb{F}_{p})\to H^{1}(G,\mathbb{F}_{p})\to\mathrm{Hom}_{\mathrm{Gal}(K(p^{\infty})/K(p))}(A,\mathbb{F}_{p})$
$\to H^{2}(\mathrm{Gal}(K(p^{\infty})/K(p)),\mathbb{F}_{p})\to H^{2}(G,\mathbb{F}_{p}).$
Here we note that
$H^{1}(\mathrm{Gal}(\Omega/K(p^{\infty})),\mathbb{F}_{p})^{\mathrm{Gal}(K(p^{\infty})/K(p))}$ $=\mathrm{Hom}_{\mathrm{Gal}(K(p^{\infty})/K(p))}(A,\mathbb{F}_{p})$
$=\oplus_{\boldsymbol{m}\in I_{0}\setminus\{(p,1)\}}\mathrm{Hom}_{\mathrm{Gal}(K(p^{\infty})/K(p))}(A^{\boldsymbol{m}},\mathbb{F}_{p}).$
To prove the assertion of (1), it suffices to compute the dimension of each eigenspace of this cohomology group. First, we compute the dimension of each eigenspace of $H^{1}(G,\mathbb{F}_{p})$. The Kummer exact sequence $1\to\mu_{p}\to\mathbb{G}_{m}\to\mathbb{G}_{m}\to 1$ gives
$1\to O_{K(p)}[1/p]^{\times}/p\to H^{1}(G,\mu_{p})\to H^{1}_{\mathrm{\acute{e}t}}(O_{K(p)}[1/p],\mathbb{G}_{m})[p]\to 0$
@@
[…]
@@
By decomposing the $p$-unit group into the product of the torsion-part and the free-part, it follows that the dimension of $\chi^{\boldsymbol{m}}$-component of $H^{1}(G,\mathbb{F}_{p})$ is at most one if $\boldsymbol{m}\in I_{0}\setminus\{(1,p),(p,1),(p-1,p-1)\}$ and otherwise at most two. By counting argument, all these inequalities are equalities. Since the action of $\Delta$ on $H^{i}(\mathrm{Gal}(K(p^{\infty})/K(p)),\mathbb{F}_{p})$ ($i=1,2$) is trivial, (1) follows from the five-term exact sequence.
By our assumption, there exists a unique prime of $K(p^{\infty})$ above $\mathfrak{p}$ which we also denote by the same letter $\mathfrak{p}$. In other words, we have an isomorphism
$\mathrm{Gal}(K(p^{\infty})_{\mathfrak{p}}/K(p)_{\mathfrak{p}})\cong\mathrm{Gal}(K(p^{\infty})/K(p)).$
@@
$\mathrm{ann}_{\mathbb{Z}_{p}[[\mathrm{Gal}(K(p^{\infty})_{\mathfrak{p}}/K(p)_{\mathfrak{p}})]]}(\mathbb{Z}_{p}(1))=\mathrm{ann}_{\Lambda}(\mathbb{Z}_{p}(1)).$
@@
Hence to prove (2), it suffices to show that the decomposition group of $A^{(1,p)}$ at $\mathfrak{p}$ coincides with the whole $A^{(1,p)}$. First, note that
$\mathrm{Hom}_{\Delta}(A^{(1,p)},\mathbb{F}_{p}(1))\cong\mathrm{Hom}_{\Delta}(\mathrm{Gal}(\Omega/K(p)),\mathbb{F}_{p}(1))\cong O_{K}[1/p]^{\times}/p$
@@
is generated by images of $\pi$ and $\bar{\pi}$. Hence the assertion is equivalent to saying that the images of $\pi$ and $\bar{\pi}$ in $K_{\mathfrak{p}}^{\times}/(K_{\mathfrak{p}}^{\times})^{p}$ is still two-dimensional. To prove the latter assertion, it suffices to show that the projection of $\bar{\pi}\in O_{K_{\mathfrak{p}}}^{\times}$ to the group of principal units $1+\pi O_{K_{\mathfrak{p}}}\cong\mathbb{Z}_{p}$ is a generator. Since the Frobenius element $\mathrm{Frob}_{\bar{\mathfrak{p}}}\in\mathrm{Gal}(K(\mathfrak{p}^{\infty})/K)\cong O_{K_{\mathfrak{p}}}^{\times}/O_{K}^{\times}$ at $\bar{\mathfrak{p}}$ is a generator by our assumption on the number of primes above $p$, we obtain the desired assertion. ∎
@@
In fact, we can completely determine the $\Lambda$-module structure of $A$ and the structure of the pro-$p$ group $\mathrm{Gal}(\Omega/K(p^{\infty}))$, though they are not necessary to prove Theorem 2.12:
Remark 4.3.
(1) We keep the same assumption as Lemma 4.2. It follows that there is a surjection
$f\colon\Lambda^{[K(p):K]-1}\oplus\mathrm{ann}_{\Lambda}(\mathbb{Z}_{p}(1))\to\mathrm{Gal}(\Omega/K(p^{\infty}))^{\mathrm{ab}}$
between $\Lambda$-modules. We claim that $f$ is an isomorphism. Let $\mathrm{Frac}(\Lambda)$ denote the field of fractions of $\Lambda$. Then we have
$\dim_{\mathrm{Frac}(\Omega)}\mathrm{ann}_{\Lambda}(\mathbb{Z}_{p}(1))\otimes_{\Lambda}\mathrm{Frac}(\Omega)=1.$
Moreover, by [NQD84, Corollaire 2.7], we have
$\dim_{\mathrm{Frac}(\Omega)}\mathrm{Gal}(\Omega/K(p^{\infty}))^{\mathrm{ab}}\otimes_{\Lambda}\mathrm{Frac}(\Omega)=[K(p):K].$
Hence it follows that $\ker(f)$ is a torsion $\Lambda$-module. However, the $\Lambda$-module $\mathrm{ann}_{\Lambda}(\mathbb{Z}_{p}(1))$ is torsion-free since it is an ideal of $\Lambda$. Hence the kernel of $f$ is trivial, as desired.
@@
(2) As a consequence of the previous remark, the $\Lambda$-module $\mathrm{Gal}(\Omega/K(p^{\infty}))^{\mathrm{ab}}$ is $\mathbb{Z}_{p}$-torsion free. Moreover, we have $H^{2}(\mathrm{Gal}(\Omega/K(p^{\infty})),\mathbb{Q}_{p}/\mathbb{Z}_{p})=0$, i.e. the weak Leopoldt conjecture holds for the $\mathbb{Z}_{p}^{2}$-extension $K(p^{\infty})/K(p)$, since it contains the cyclotomic $\mathbb{Z}_{p}$-extension by [NQD84, Theoreme 2.2]. Now, by considering a long exact sequence associated to
$0\to\mathbb{F}_{p}\to\mathbb{Q}_{p}/\mathbb{Z}_{p}\xrightarrow{\times p}\mathbb{Q}_{p}/\mathbb{Z}_{p}\to 0,$
it follows that $H^{2}(\mathrm{Gal}(\Omega/K(p^{\infty})),\mathbb{F}_{p})=0$. In other words, the pro-$p$ group $\mathrm{Gal}(\Omega/K(p^{\infty}))$ is a free pro-$p$ group. The proof of Theorem 2.12 below will provide an explicit basis of this free pro-$p$ group.
Now we define elements $\sigma_{\boldsymbol{m}}\in F^{\boldsymbol{m}}G_{K}$ for $\boldsymbol{m}\in I_{0}$ as follows:
@@
Construction. For $\boldsymbol{m}\in I_{0}$, we choose an element $t_{\boldsymbol{m}}\in\mathrm{Gal}(\Omega/K(p^{\infty}))$, as follows:
@@
  • If $\boldsymbol{m}\in I_{0}\setminus\{(p,1),(1,p),(p-1,p-1)\}$, we choose a lift $t_{\boldsymbol{m}}\in\mathrm{Gal}(\Omega/K(p^{\infty}))$ of a generator of $A^{\boldsymbol{m}}$ as $\mathbb{Z}_{p}[[\mathrm{Gal}(K(p^{\infty})/K(p))]]$-module.
@@
Let $n\geq 2$ be an arbitrary integer such that $n\equiv 1\bmod p-1$. Then we have $\kappa_{(1,n)}(t_{(p,1)})=0$. Similarly, for such $n$, we have $\kappa_{(n,1)}(t_{(1,p)})=0$.
@@
$[\Pi(2),\Pi(2)]=[\Pi(1,1),\Pi(1,1)]\subset\Pi(2,2)$
and $[x_{1},[x_{2},z]]\in\Pi(2,2)$, we have $s(t_{(p,1)})(x_{2})x_{2}^{-1}\in\Pi(2,2)$ as desired.
@@
Observe that the action of $x_{2}$ sends $[w_{i},w_{j}]\in\Pi(2,0)/\Pi(3,0)$ to
$[w_{i+1}w_{i},w_{j+1}w_{j}]=[w_{i+1},w_{j+1}]+[w_{i+1},w_{j}]+[w_{i},w_{j+1}]+[w_{i},w_{j}].$
@@
Next we prove the assertion of (2). For $\boldsymbol{m}\in I_{0}\setminus\{(1,p),(p,1),(p-1,p-1)\}$, the assertion immediately follows since we have $\kappa_{\boldsymbol{m}}(t_{\boldsymbol{m}})=\kappa_{\boldsymbol{m}}(\sigma_{\boldsymbol{m}})$ and the element $t_{\boldsymbol{m}}$ generates $A^{\boldsymbol{m}}$ as a $\Lambda$-module. Since the $\boldsymbol{m}$-th Soulé character is nontrivial by Theorem 2.6 (2), the image $\kappa_{\boldsymbol{m}}(F^{1}G)$ is nonzero.
Assume that $\boldsymbol{m}=(p-1,p-1)$. Since $A^{(p-1,p-1)}=A^{\Delta}$, the $\Delta$-invariant part of $A$, it follows that $A^{(p-1,p-1)}$ is isomorphic to $\mathrm{Gal}(L/K_{\infty})^{{\rm ab}}$, where $L$ is the maximal pro-$p$ extension of $K$ unramified outside $p$. Since $K$ has class number one, it follows by [NSW08, (10.7.13) Theorem] that $\mathrm{Gal}(L/K)$ is a free pro-$p$ group of rank two on the set $\{\gamma_{1},\gamma_{2}\}$. Hence $\mathrm{Gal}(L/K_{\infty})^{\rm{ab}}$, which is just the maximal abelian quotient of the commutator subgroup of a free pro-$p$ group of rank two, is a free $\Lambda$-module of rank one generated by $[\gamma_{1},\gamma_{2}]$ by [Iha86, Theorem 2]. Hence the image of $\kappa_{(p-1,p-1)}(\sigma_{(p-1,p-1)})$ generates $\kappa_{(p-1,p-1)}(A)$, which is nonzero by Theorem 2.6 (2). Finally, the case where $\boldsymbol{m}=(p,1)$ or $(1,p)$ follows from Lemma 4.5 and Theorem 2.6 (2). ∎
@@
Since $A$ is abelian, the elements $\epsilon_{1,m}$ and $\epsilon_{2,m}$ act on $A$ as idempotents for every $m\geq 0$. Moreover, the action of $\epsilon_{i,m}$ commutes with the conjugation by $\gamma_{1}$ and $\gamma_{2}$. Hence the assertion follows from the construction of $\sigma_{\boldsymbol{m}}$ and $g_{\boldsymbol{m}}$. ∎
We now show that $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ satisfy the assumption of Conjecture 2.8:

### v3

@@ 4.1. Construction of elements
In this subsection, we construct elements $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ satisfying the assumption of Conjecture 2.3 in such a way that can be regarded as a two-variable variant of Sharifi’s construction in [Sha02, 2]. We also construct auxiliary elements $\{g_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$, which are similar to $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ but much easier to handle. These auxiliary elements play an important role when proving that $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ strongly generate $F^{1}G$, see the proof of Theorem 2.14 in Section 4.2 for more details.
First, we lift generators of $\operatorname{Gal}(K(p^{\infty})/K)\cong\Delta\times\mathbb{Z}_{p}^{2}$ to $\operatorname{Gal}(\Omega/K)$, as follows: Let us denote the maximal pro-$p$ subextension of $K(p^{\infty})/K$ by $K_{\infty}/K$. Then it is a $\mathbb{Z}_{p}^{2}$-extension. The upper exact sequence in the following commutative diagram
@@
$\scriptstyle\sim$
splits since $\operatorname{Gal}(\Omega/K(p^{\infty}))$ is a pro-$p$ group and $\Delta$ is a prime-to-$p$ group by Schur-Zassenhaus [Ser97, 5.9, Corollary 1]. We fix a section $r\colon\operatorname{Gal}(K(p)/K)\to\operatorname{Gal}(\Omega/K_{\infty})$ and identify $\operatorname{Gal}(\Omega/K)$ with the semi-direct product $G\rtimes\Delta$. Then $\Delta$ acts on $\operatorname{Gal}(\Omega/K(p^{\infty}))$ through this section by conjugation, and
$1\to\operatorname{Gal}(\Omega/K(p^{\infty}))\to G\to\operatorname{Gal}(K(p^{\infty})/K(p))\to 1$
is an exact sequence of pro-$p$ groups with $\Delta$-action, noting that the action of $\Delta$ on $\operatorname{Gal}(K(p^{\infty})/K(p))$ is trivial. Then the sequence
$1\to\operatorname{Gal}(\Omega/K(p^{\infty}))^{\Delta}\to G^{\Delta}\to\operatorname{Gal}(K(p^{\infty})/K(p))\to 1$
@@
$G^{\Delta}\to\operatorname{Gal}(K(p^{\infty})/K(p))$
is surjective by applying Lemma 4.2 below to the case of $m=0$ below.
Let $\gamma_{1}$ (resp. $\gamma_{2}$) be an element of the $\Delta$-invariant subgroup $G^{\Delta}$ which restricts to a generator of $\operatorname{Gal}(K(p^{\infty})/K(\bar{\mathfrak{p}}^{\infty}\mathfrak{p}))\cong\mathbb{Z}_{p}$ (resp. $\operatorname{Gal}(K(p^{\infty})/K(\mathfrak{p}^{\infty}\bar{\mathfrak{p}}))\cong\mathbb{Z}_{p}$). Moreover, we fix a generator $\delta\in\mathbb{F}_{p}^{\times}$ and let $\delta_{1}$ (resp. $\delta_{2}$) be an element of $\operatorname{Gal}(\Omega/K)$ defined to be the image of the image of $(\delta,1)$ (resp. $(1,\delta)$) in $\Delta$ through the homomorphism
$(\mathbb{F}_{p}^{\times}\times\mathbb{F}_{p}^{\times})/\operatorname{Im}(O_{K}^{\times})\xrightarrow{(\chi_{1},\chi_{2})^{-1}}\Delta\xrightarrow{r}\operatorname{Gal}(\Omega/K),$
where the first arrow is an isomorphism induced by characters $\chi_{1}$ and $\chi_{2}$. Then $\delta_{1}$ (resp. $\delta_{2}$) is contained in $\operatorname{Gal}(\Omega/K(\bar{\mathfrak{p}}))$ (resp. $\operatorname{Gal}(\Omega/K(\bar{p}))$) and, by construction, we have the following relations:
$\delta_{1},\delta_{2}]=1,\quad[\delta_{1},\gamma_{1}]=1,\quad[\delta_{1},\gamma_{2}]=1,\quad[\delta_{2},\gamma_{1}]=1\quad\text{and}\quad[\delta_{2},\gamma_{2}]=1.\quad$
For $m\in\mathbb{Z}_{\geq 0}$ and $1\leq i\leq 2$, we define an idempotent $\epsilon_{i,m}\in\mathbb{Z}_{p}[\operatorname{Gal}(\Omega/K)]$ by
@@
For every $g\in G$, $m\in\mathbb{Z}_{\geq 0}$ and $1\leq i\leq 2$, the limit
@@
Let $A$ denote the maximal abelian quotient of $F^{1}G=\operatorname{Gal}(\Omega/K(p^{\infty}))$, which is naturally endowed with the structure of a $\Lambda\coloneqq\mathbb{Z}_{p}[[\operatorname{Gal}(K(p^{\infty})/K(p))]]$-module. In the following lemma and remark, we determine the structure of this $\Lambda$-module completely:
@@
Suppose the class number of $K(p)$ is not divisible by $p$ and there are exactly two primes of $K(p^{2})$ lying above $p$. Let $\boldsymbol{m}\in I_{0}$ and let $A^{\boldsymbol{m}}\coloneqq\epsilon_{1,m_{1}}\epsilon_{2,m_{2}}A$. Then the following assertions hold:
@@
    $A^{(p,1)}=A^{(1,p)}$ is a quotient of the annihilator ideal $\operatorname{Ann}_{\Lambda}(\mathbb{Z}_{p}(1))$ of $\mathbb{Z}_{p}(1)$.
@@
$0$ $\to H^{1}(\operatorname{Gal}(K(p^{\infty})/K(p)),\mathbb{F}_{p})\to H^{1}(G,\mathbb{F}_{p})\to\operatorname{Hom}_{\operatorname{Gal}(K(p^{\infty})/K(p))}(A,\mathbb{F}_{p})$
$\to H^{2}(\operatorname{Gal}(K(p^{\infty})/K(p)),\mathbb{F}_{p})\to H^{2}(G,\mathbb{F}_{p}).$
Note that $\operatorname{Hom}_{\operatorname{Gal}(K(p^{\infty})/K(p))}(A,\mathbb{F}_{p})=\oplus_{\boldsymbol{m}\in I_{0}\setminus\{(p,1)\}}\operatorname{Hom}_{\operatorname{Gal}(K(p^{\infty})/K(p))}(A^{\boldsymbol{m}},\mathbb{F}_{p}).$ To prove the assertion of (1), it suffices to compute the dimension of each eigenspace of this cohomology group. First, we compute the dimension of each eigenspace of $H^{1}(G,\mathbb{F}_{p})$. The Kummer exact sequence $1\to\mu_{p}\to\mathbb{G}_{m}\to\mathbb{G}_{m}\to 1$ gives
$1\to O_{K(p)}[1/p]^{\times}/p\to H^{1}(G,\mu_{p})\to H^{1}_{\mathrm{\acute{e}t}}(O_{K(p)}[1/p],\mathbb{G}_{m})[p]\to 0.$
@@
Since there are only two primes of $K(p)$ lying above $p$ by assumption, the dimension of the left hand side is equal to $[K(p):K]+2$ by Dirichlet’s unit theorem. Moreover, by [NSW08, (8.7.2) Proposition], there is an isomorphism of $\mathbb{Q}[\Delta]$-modules
@@
By decomposing the $p$-unit group into the product of the torsion-part and the free-part, it follows that the dimension of $\chi^{\boldsymbol{m}}$-component of $H^{1}(G,\mathbb{F}_{p})$ is at most one if $\boldsymbol{m}\in I_{0}\setminus\{(1,p),(p,1),(p-1,p-1)\}$ and otherwise at most two. By a counting argument, all these inequalities are equalities. Since $\Delta$ acts trivially on $H^{i}(\operatorname{Gal}(K(p^{\infty})/K(p)),\mathbb{F}_{p})$ ($i=1,2$), (1) follows from the five-term exact sequence.
By our assumption, there exists a unique prime of $K(p^{\infty})$ above $\mathfrak{p}$ which we denote by the same letter $\mathfrak{p}$ (see Remark 2.14.1). Hence we have an isomorphism
$\operatorname{Gal}(K(p^{\infty})_{\mathfrak{p}}/K(p)_{\mathfrak{p}})\xrightarrow{\sim}\operatorname{Gal}(K(p^{\infty})/K(p)).$
@@
$\operatorname{Ann}_{\mathbb{Z}_{p}[[\operatorname{Gal}(K(p^{\infty})_{\mathfrak{p}}/K(p)_{\mathfrak{p}})]]}(\mathbb{Z}_{p}(1))=\operatorname{Ann}_{\Lambda}(\mathbb{Z}_{p}(1)).$
@@
Hence to prove (2), it suffices to show that $A^{(1,p)}$ coincides with its decomposition subgroup at $\mathfrak{p}$. First, note that
$\operatorname{Hom}_{\Delta}(A^{(1,p)},\mathbb{F}_{p}(1))\cong\operatorname{Hom}_{\Delta}(\operatorname{Gal}(\Omega/K(p)),\mathbb{F}_{p}(1))\cong O_{K}[1/p]^{\times}/p\cong\mathbb{F}_{p}^{2}$
@@
is generated by $\pi$ and $\bar{\pi}$. Hence the assertion is equivalent to saying that the images of $\pi$ and $\bar{\pi}$ in $K_{\mathfrak{p}}^{\times}/(K_{\mathfrak{p}}^{\times})^{p}$ still span a two-dimensional subspace. This assertion is then equivalent to saying that the projection of $\bar{\pi}\in O_{K_{\mathfrak{p}}}^{\times}$ to the group of principal units $1+\mathfrak{p}O_{K_{\mathfrak{p}}}\cong\mathbb{Z}_{p}$ is a generator. However, this follows since the Frobenius element at $\bar{\mathfrak{p}}$ in $\operatorname{Gal}(K(\mathfrak{p})/K)$, which coincides with $\bar{\pi}$ under $\operatorname{Gal}(K(\mathfrak{p})/K)\cong(O_{K_{\mathfrak{p}}}/\mathfrak{p})^{\times}/O_{K}^{\times}$, is a generator by our assumption on the number of primes above $p$. ∎
@@
In fact, we can determine the $\Lambda$-module structure of $A$ and the structure of the pro-$p$ group $\operatorname{Gal}(\Omega/K(p^{\infty}))$ completely, though they are not necessary to prove Theorem 2.14. Here we record the precise statement and provide its proof for the interested reader:
Proposition 4.4.
Suppose the class number of $K(p)$ is not divisible by $p$ and there are exactly two primes of $K(p^{2})$ above $p$. The following assertions hold.
  1. (1)
    There exists an isomorphism
    $\Lambda^{[K(p):K]-1}\oplus\operatorname{Ann}_{\Lambda}(\mathbb{Z}_{p}(1))\xrightarrow{\sim}A$
    between $\Lambda$-modules.
  2. (2)
    The group $\operatorname{Gal}(\Omega/K(p^{\infty}))$ is a free pro-$p$ group (of countably infinite rank).
Proof.
(1) The proof of Lemma 4.3 shows that there exists a surjective homomorphism
$f\colon\Lambda^{[K(p):K]-1}\oplus\operatorname{Ann}_{\Lambda}(\mathbb{Z}_{p}(1))\to A$
between $\Lambda$-modules. Moreover, by [NQD84, Corollaire 2.7], we have
$\dim_{\mathrm{Frac}(\Omega)}A\otimes_{\Lambda}\mathrm{Frac}(\Omega)=[K(p):K].$
This forces the kernel of $f$ to be a torsion $\Lambda$-module. However, since the left-hand side is obviously torsion-free, the assertion follows.
@@
(2) The assertion (1) implies that $A=\operatorname{Gal}(\Omega/K(p^{\infty}))^{\mathrm{ab}}$ is $\mathbb{Z}_{p}$-torsion free. Moreover, we have $H^{2}(\operatorname{Gal}(\Omega/K(p^{\infty})),\mathbb{Q}_{p}/\mathbb{Z}_{p})=0$, i.e. the weak Leopoldt conjecture holds for the $\mathbb{Z}_{p}^{2}$-extension $K(p^{\infty})/K(p)$ by [NQD84, Theoreme 2.2], since it contains the cyclotomic $\mathbb{Z}_{p}$-extension. Now, by considering a long exact sequence associated to
$0\to\mathbb{F}_{p}\to\mathbb{Q}_{p}/\mathbb{Z}_{p}\xrightarrow{p}\mathbb{Q}_{p}/\mathbb{Z}_{p}\to 0,$
we obtain $H^{2}(\operatorname{Gal}(\Omega/K(p^{\infty})),\mathbb{F}_{p})=0$. This concludes the proof. ∎
Now we construct elements $\sigma_{\boldsymbol{m}}\in F^{\boldsymbol{m}}G_{K}$ for $\boldsymbol{m}\in I_{0}$. In the rest of this section, since our construction relies on Lemma 4.3, we suppose $p$ satisfies the following two assumptions of Theorem 2.14:
(1) The class number of $K(p)$ is not divisible by $p$. (2) There are exactly two primes of $K(p^{2})$ above $p$.
@@
Construction. For $\boldsymbol{m}\in I_{0}$, we choose an element $t_{\boldsymbol{m}}\in\operatorname{Gal}(\Omega/K(p^{\infty}))$, as follows:
@@
  • If $\boldsymbol{m}\in I_{0}\setminus\{(p,1),(1,p),(p-1,p-1)\}$, we choose a lift $t_{\boldsymbol{m}}\in\operatorname{Gal}(\Omega/K(p^{\infty}))$ of a generator of $A^{\boldsymbol{m}}$ as $\mathbb{Z}_{p}[[\operatorname{Gal}(K(p^{\infty})/K(p))]]$-module.
@@
Let $n\geq 2$ be an integer such that $n\equiv 1\bmod p-1$. Then we have $\kappa_{(1,n)}(t_{(p,1)})=0$. Similarly, for such $n$, we have $\kappa_{(n,1)}(t_{(1,p)})=0$.
@@
$\Pi(2),\Pi(2)]=[\Pi(1,1),\Pi(1,1)]\subset\Pi(2,2)$
and $x_{1},[x_{2},z]]\in\Pi(2,2)$, we have $s(t_{(p,1)})(x_{2})x_{2}^{-1}\in\Pi(2,2)$ as desired.
@@
Observe that the action of $x_{2}$ sends $w_{i},w_{j}]\in\Pi(2,0)/\Pi(3,0)$ to
$w_{i+1}w_{i},w_{j+1}w_{j}]=[w_{i+1},w_{j+1}]+[w_{i+1},w_{j}]+[w_{i},w_{j+1}]+[w_{i},w_{j}].$
@@
Next we prove the assertion of (2). For $\boldsymbol{m}\in I_{0}\setminus\{(1,p),(p,1),(p-1,p-1)\}$, the assertion immediately follows since we have $\kappa_{\boldsymbol{m}}(t_{\boldsymbol{m}})=\kappa_{\boldsymbol{m}}(\sigma_{\boldsymbol{m}})$ and the element $t_{\boldsymbol{m}}$ generates $A^{\boldsymbol{m}}$ as a $\Lambda$-module. Since the character $\kappa_{\boldsymbol{m}}$ is nontrivial by Lemma 4.1, the image $\kappa_{\boldsymbol{m}}(F^{1}G)$ is nonzero.
Assume that $\boldsymbol{m}=(p-1,p-1)$. Since $A^{(p-1,p-1)}=A^{\Delta}$, the $\Delta$-invariant part of $A$, it follows that $A^{(p-1,p-1)}$ is isomorphic to $\operatorname{Gal}(L/K_{\infty})^{{\rm ab}}$, where $L$ is the maximal pro-$p$ extension of $K$ unramified outside $p$. Since $K$ has class number one, it follows by [NSW08, (10.7.13) Theorem] that $\operatorname{Gal}(L/K)$ is a free pro-$p$ group of rank two on the set $\{\gamma_{1},\gamma_{2}\}$. Hence $\operatorname{Gal}(L/K_{\infty})^{\rm{ab}}$, which is just the maximal abelian quotient of the commutator subgroup of a free pro-$p$ group of rank two, is a free $\Lambda$-module of rank one generated by $\gamma_{1},\gamma_{2}]$ by [Iha86, Theorem 2]. Hence the image of $\kappa_{(p-1,p-1)}(\sigma_{(p-1,p-1)})$ generates $\kappa_{(p-1,p-1)}(A)$, and it is nontrivial by Lemma 4.1. Finally, the case where $\boldsymbol{m}=(p,1)$ or $(1,p)$ follows from Lemma 4.1 and Lemma 4.6. ∎
@@
Since $A$ is abelian, the elements $\epsilon_{1,m}$ and $\epsilon_{2,m}$ act on $A$ as idempotents for every $m$. Moreover, the action of $\epsilon_{i,m}$ commutes with the conjugation by $\gamma_{1}$ and $\gamma_{2}$. Hence the assertion follows from the construction of $\sigma_{\boldsymbol{m}}$ and $g_{\boldsymbol{m}}$. ∎
We now show that $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ satisfy the assumption of our conjecture 2.3: